# LaVA-Man: Learning Visual Action Representations for Robot Manipulation

### v1

@@ 2 Related work
Visual pre-training for robotics. Visual pre-training can be applied to robotic manipulation to enhance the generalisation ability [7, 37]. They rely on various pre-training methods, such as contrastive learning [38, 39], MAE [7, 40, 41], and other perception tasks [42, 43]. For instance, the approach in [37] pre-trains a visual backbone using classic tasks such as image classification and object detection. VIP [44], R3M [45], MVP [40, 41], and MCR [46] aim to learn implicit features for robotic tasks by training on video data, while Thiea [47] and SUSIE [48] try to distil knowledge from pre-trained vision foundation models. SUGAR [43] and 3D-MVP [49] extend the pre-training tasks to the 3D domain. To further improve the semantic understanding, Voltron [7] and MPI [42] combine video-based tasks with text conditioning, enabling generalisation to diverse object manipulation tasks. However, video data often contains temporal redundancy, which may allow models to exploit smooth temporal transitions rather than learning meaningful representations [50]. On the other hand, we design a new pretext paradigm to directly train on robot image pairs with textual instructions via asymmetric masking [34, 36].
@@ 3 Method
Given a visual observation ${\mathbf{o}}_{s}$ and associated language instruction ${\mathbf{l}}_{s\to f}$, our goal is to learn a policy that predicts a robot action ${\mathbf{a}}_{s\to f}$ that leads to a goal observation ${\mathbf{o}}_{f}$. We formulate this problem into two stages: 1) Visual-action representation learning: We first train a model $f_{\theta}$ that outputs the predicted goal image, $\hat{{\mathbf{o}}}_{f}$, from ${\mathbf{o}}_{s}$ and ${\mathbf{l}}_{s\to f}$. This self-supervised task encourages the model to align visual and textual modalities, capturing how the instruction transforms the scene. 2) Robot action prediction: We then fine-tune the model by attaching an action prediction head that maps the learned representation to a robot action $\hat{{\mathbf{a}}}_{s\to f}$. This action may take various forms depending on the task setting, e.g., $\hat{{\mathbf{a}}}_{s\to f}\in SE(2)$ for tabletop manipulation or $\hat{{\mathbf{a}}}_{s\to f}\in\mathbb{R}^{9}$ for joint-angle visuomotor control.
@@ 3.1 Learning visual-action representations
We propose a self-supervised approach of asymmetric masking to learn visual-action representations using the pretext task, which aims to reconstruct a partially masked goal image given ${\mathbf{o}}_{s}$ and ${\mathbf{l}}_{s\to f}$. Fig. 1 shows LaVA-Man’s architecture and the proposed pretext task, goal image prediction.
@@ 4 The Omni-Object Pick-and-Place dataset
Data annotation. OOPP is built upon the previous benchmarks [8, 9] in the Pybullet Gym environment [11]. We adopt the objects from the OmniObject3D dataset [10], filtering out the objects unsuitable for tabletop manipulation by retaining only objects with diameters between 4cm and 40cm. During simulation, we allow the robot to perform automatic manipulation with a pre-defined placeholder for collecting action annotations. These placeholders are then replaced with real objects. To ensure spatial diversity and prevent overcrowding, we adopt a KDTree-based spatial partitioning strategy inspired by [1], subdividing scenes into feasible regions for object placement. Based on the rich language description for each object provided by the original OmniObjects3D dataset [10], we generate diverse language instructions that are beyond just using object names, offering more natural and varied descriptions of object appearances as in Fig. 2.
@@ 5.1 Setup
Table 3: Results on Referring Expression Grounding [7] and Franka Kitchen [54]. Our method outperforms self-supervised methods and is comparable to leading supervised methods. The results of existing methods are from [42] and [7]. Minimum, Medium, Maximum in (a) denote the scenarios categorised according to the level of clutter.
@@
Figure 3: Real-robot experiments. We report the   perceptual score and   physical score, which indicate the success rates in affordance detection and physical robotic tasks, respectively. The tasks of Opening drawer and Pushing piles are not included in the training demonstrations.
@@ 5.2 Evaluation
Referring expression grounding results in Tab. 3 (a), demonstrate that our method achieves better or comparable results, even compared with supervised methods [42, 43]. This validates that the goal-image prediction pretext task can learn the association between the target object and the corresponding language instructions, resulting in accurate object localisation.
@@ 7 Limitations
Articulated objects. In real-world scenarios, some objects are composed of multiple rigid parts and interactable. However, our visual pre-training primarily focuses on rigid object manipulation and does not explicitly model any dynamics or interactions. Including articulated objects with several new tasks in our OOPP dataset can potentially help capture these complex interactions.
@@ 1 Introduction
  • We invite readers to watch the demo video about the proposed pretext task (goal-image prediction), Omni-Object Pick-and-Place (OOPP) dataset, and real and simulated robot manipulation.
@@ 2 Method details
Backbone architecture. We adopt ViT-Base [61] with a patch size of 16 as our backbone. The original Vit-Base architecture consists of 12 attention blocks with an embedding dimension of 768, and its the decoder typically comprises 8 attention blocks with an embedding dimension of 512. For a fair comparison with other methods using ViT-Base as a backbone, we include 6 attention blocks and 6 bi-directional attention blocks in our encoder. In the pre-training stage, our goal-image prediction head is a one-layer fully connected network.
@@ 3.1 Ravens
Overview. Ravens [8] is a simulated benchmark to evaluate tabletop pick-and-place robot manipulation tasks. We use PyBullet OpenAI Gym [11] based on the configuration described in CLIPort [1]. We chose 8 language-conditioned tasks for the experiment, as shown in 7, including Packing seen or unseen Google objects sequence, Packing seen or unseen Google objects group, Put block in the bowl, Stack blocking pyramid, Towers of Hanoi, Packing boxes pairs, Assembling kits, and Separating piles. Details of each task, including the train and test split of objects, and the language instruction template, can be referred to  [1]. Note that we did not split the tasks according to seen or unseen colours, as all the methods have already “seen” all the colours in their pre-train models or the unsupervised pertaining phase. Therefore, we combine both “seen” and “unseen” splits of colours into a single task and the scores just reflect all model’s perception ability on all the colours.
Evaluation details. We evaluated the capability of the proposed methods on multi-task experiments in the benchmark. Specifically, we trained the model using 1,000 demonstrations drawn from all task categories and assessed performance on another 100 test demonstrations per task. Since prior pre-training methods [42, 7] have not been evaluated on this benchmark, we reimplemented their models using the original codebases and pre-trained them on the same pre-training data as our approach. When adapting to the downstream tasks, these pre-training methods were also equipped with the affordance head described in Section 2. Subsequently, all baselines, including the Pick-and-Place baselines, were fully fine-tuned on the same set of downstream demonstrations, following the evaluation protocol outlined in [1], to ensure a fair comparison.
@@ 3.2 Franka Kitchen
Overview. Franka Kitchen [54] is a well-established benchmark for evaluating the efficacy of visual representations in facilitating the learning of visuomotor control policies from limited demonstrations. This benchmark comprises five distinct visuomotor control tasks, as shown in the Fig. 8, each captured from two camera viewpoints.
@@ 3.3 Referring Expression Grounding
Overview. The goal of this task is to predict a bounding box of an object in a cluttered scene based on the nature language expression. This task offers the evaluation of language-conditioned scene understanding and object recognising ability, which serves as an important prerequisite for language-based robot manipulation. The benchmark is based on OCID-Ref Dataset [62], which provides representatives scenes in robotises settings. The benchmark also provides splits based on the clutter level.
@@
Evaluation details. We use a shallow MLP as detection head to regress the bounding box directly. We use the evaluation codebase provided by [7]. Similar to Franka Kitche, we report the results stated in the paper for each baseline [7, 42, 45, 40, 43]. The evaluation metrics are the average precision at 0.25 IoU under each clutter level.
@@ 3.4 Real robot experiments
Evaluation details. Fig. 10 shows our real-robot environment and objects. We trained both our model and CLIPort [1] on our manually collected training demos. We utilise a 6-degree-of-freedom (6-DoF) UR5 robotic arm, Robotiq 2F-85 two-finger gripper, and an Intel RealSense RGB-D camera for our real-world experiments. We capture the top-down RGB observation which covers the workspace of 60 cm $\times$ 30 cm, and the image from the camera is resized to 320 $\times$ 160 pixels.
Task details. Here we show the language template, variable factors and the success condition for each real robot task in Tab. 4. Images for each task could be referred in the main paper.
@@
Press button Press the {color} button Button color, button location Correct button touched
@@ 4 Additional details
Figure 12: Examples of masked image training. We show the qualitative examples of our masked auto-encoders. The text descriptions are shown below each sample.
@@
Dataset details. We build our Omni-Object Pick-and-Place (OOPP) dataset upon the previous benchmarsk [1, 9] in the PyBullet Gym enviornment [11]. We manually selected 180 real-scanned object classes from the OmniObject3D dataset [10], focusing on those suitable for robotic manipulation, resulting in a total of 3,200 distinct instances. For each object, we reduced the mesh resolution to 20K faces to enable efficient rendering in simulation. Additionally, we filtered out objects that were either too large or too small i.e., with dimensions less than 4cm or greater than 40cm). Our dataset includes full robot manipulation episodes, comprising annotated robot actions, language instructions, and simulation-generated rewards. We use 160 object classes for training. For evaluating intra-class generalisation, we hold out a subset of instances from 20 categories included in the training set. For inter-class generalisation, we reserve 20 object categories that are entirely unseen during training. Fig. 13 illustrates the first 50 examples across different object classes, while Fig. 14 highlights examples of intra-class variation. The held-out categories for inter-class generalisation are sampled from four high-level semantic groups, as detailed in Table 5.
Table 5: Semantic group division of seen and unseen classesWe divided all objects into four semantic groups. We selected unseen classes from each group in proportion to the number of object classes it contains, ensuring an even distribution of unseen categories.
@@
dumpling, mooncake, rice_cake, puff, water_chestnut, mushroom, egg,
@@
peach, apple, pear, potato, zongzi, pomegranate, sweet_potato, onion,
@@
belt, dish, flash_light, canned_beverage, fork, cup
glasses_case, bowl, tape_measure, speaker, laundry_detergent, teapot
@@
picnic_basket, hammer, shoe, hat, laptop, vase, ornaments, spanner, book,
@@
toy_boat, toy_plant, toy_car, toy_bus, toy_plane, timer, whistle,
table_tennis_bat, toy_motorcycle, drum, remote_control, toy_animals,
garage_kit, china, chess, Chinese_chess, rubik_cube, dinosaur, doll
@@
hairpin, lotus_root, house (model), plant, dumbbell, package
@@
Figure 13: Examples of objects from different classes.
Figure 14: Examples of intra-class variance.

### v2

@@ 2 Related work
Visual pre-training for robotics. Visual pre-training can be applied to robotic manipulation to enhance the generalisation ability [7, 37]. They rely on various pre-training methods, such as contrastive learning [38, 39], MAE [7, 40, 41], and other perception tasks [42, 43]. For instance, the approach in [37] pre-trains a visual backbone using classic tasks such as image classification and object detection. VIP [44], R3M [45], MVP [40, 41], and MCR [46] aim to learn implicit features for robotic tasks by training on video data, while Thiea [47] and SUSIE [48] try to distil knowledge from pre-trained vision foundation models. SUGAR [43] and 3D-MVP [49] extend the pre-training tasks to the 3D domain. To further improve the semantic understanding, Voltron [7] and MPI [42] combine video-based tasks with text conditioning, enabling generalisation to diverse object manipulation tasks. However, video data often contains temporal redundancy, which may allow models to exploit smooth temporal transitions rather than learning meaningful representations [50]. On the other hand, we design a new pretext paradigm to directly predict goal image with textual instructions via asymmetric masking [34, 36].
@@ 3 Method
Given a visual observation ${\mathbf{o}}_{s}$ and associated language instruction ${\mathbf{l}}_{s\to f}$, our goal is to learn a policy that predicts a robot action ${\mathbf{a}}_{s\to f}$ that leads to a goal observation ${\mathbf{o}}_{f}$. We formulate this problem into two stages: 1) Visual-action representation learning: We first train a model $f_{\theta}$ that outputs the predicted goal image, $\hat{{\mathbf{o}}}_{f}$, from ${\mathbf{o}}_{s}$ and ${\mathbf{l}}_{s\to f}$. This self-supervised task encourages the model to align visual and textual modalities, capturing how the instruction transforms the scene. 2) Robot action prediction: We then fine-tune the model by attaching an action prediction head that maps the learned representation to a robot action $\hat{{\mathbf{a}}}_{s\to f}$. This action may take various forms depending on the task setting, e.g., $\hat{{\mathbf{a}}}_{s\to f}\in SE(2)$ for tabletop manipulation or $\hat{{\mathbf{a}}}_{s\to f}\in\mathbb{R}^{9}$ for joint angle control.
@@ 3.1 Learning visual-action representations
We propose a self-supervised approach of asymmetric masking to learn visual-action representations using the pretext task, which aims to reconstruct a partially masked goal image given ${\mathbf{o}}_{s}$ and ${\mathbf{l}}_{s\to f}$. Fig. 1 shows LaVA-Man’s architecture and the proposed pretext task of goal image prediction.
@@ 4 The Omni-Object Pick-and-Place dataset
Data annotation. OOPP is built upon the previous benchmarks [8, 9] in the Pybullet Gym environment [11]. We adopt the real-scanned objects from the OmniObject3D dataset [10], filtering out the objects unsuitable for tabletop manipulation by retaining only objects with diameters between 4cm and 40cm. During simulation, we allow the robot to perform automatic manipulation with a pre-defined placeholder for collecting action annotations. These placeholders are then replaced with real objects. To ensure spatial diversity and prevent overcrowding, we adopt a KDTree-based spatial partitioning strategy inspired by [1], subdividing scenes into feasible regions for object placement. Based on the rich language description for each object provided by the original OmniObjects3D dataset [10], we generate diverse language instructions that are beyond just using object names, offering more natural and varied descriptions of object appearances as in Fig. 2.
@@ 5.1 Setup
Table 3: Results on Referring Expression Grounding [7] and Franka Kitchen [54]. Our method outperforms self-supervised methods and is comparable to leading supervised methods. The results of existing methods are quoted from [42] and [7]. Minimum, Medium and Maximum in sub-table (a) denote the scenarios categorised according to the level of clutter.
@@
Figure 3: Real-robot experiments. We report the   perceptual score and   physical score, which indicate the success rates for affordance perception and physical robotic tasks, respectively. The tasks of Opening drawer and Pushing piles are not included in the training demonstrations.
@@ 5.2 Evaluation
Referring expression grounding results in Tab. 3 (a) demonstrate that our method achieves better or comparable results, even compared with supervised methods [42, 43]. This validates that the goal-image prediction pretext task can learn the association between the target object and the corresponding language instructions, resulting in accurate object localisation.
@@ 7 Limitations
Articulated objects. In real-world scenarios, some objects are composed of multiple rigid parts and interactable. However, our visual pre-training primarily focuses on rigid object manipulation and does not explicitly model object-part interactions. Including articulated objects with several new tasks in our OOPP dataset can potentially help capture these complex interactions.
@@ 1 Introduction
  • We invite readers to visit our website about the proposed pretext task (goal-image prediction), Omni-Object Pick-and-Place (OOPP) dataset, and real and simulated robot manipulation.
@@ 2 Method details
Backbone architecture. We adopt ViT-Base [61] with a patch size of 16 as our backbone. The original Vit-Base architecture consists of 12 attention blocks with an embedding dimension of 768, and its the decoder typically comprises of 8 attention blocks with an embedding dimension of 512. For a fair comparison with other methods using ViT-Base as a backbone, we include 6 attention blocks and 6 bi-directional attention blocks in our encoder. In the pre-training stage, our goal-image prediction head is a one-layer fully connected network.
@@ 3.1 Ravens
Overview. Ravens [8] is a simulated benchmark to evaluate tabletop pick-and-place robot manipulation tasks. We use PyBullet OpenAI Gym [11] based on the configuration described in CLIPort [1]. We chose 8 language-conditioned tasks for the experiment, as shown in Fig. 7, including Packing seen or unseen Google objects sequence, Packing seen or unseen Google objects group, Put block in the bowl, Stack blocking pyramid, Towers of Hanoi, Packing boxes pairs, Assembling kits, and Separating piles. For details of each task, including the train and test split of objects, and the language instruction template, please refer to [1, 8]. Note that we did not split the tasks according to seen or unseen colours, as all the methods have already “seen” all the colours in their pre-train models or the unsupervised pre-training phase. Therefore, we combine both “seen” and “unseen” splits of colours into a single task and the scores reflect the model’s perception ability on all the colours.
Evaluation details. We evaluated the capability of the proposed methods on multi-task experiments in the benchmark. Specifically, we trained the model using 1,000 demonstrations drawn from all task categories and assessed performance on another 100 test demonstrations per task. Since prior pre-training methods [7, 42] have not been evaluated on this benchmark, we reimplemented their models using the original codebases and pre-trained them on the same pre-training data as our approach. When adapting to the downstream tasks, these pre-training methods were also equipped with the affordance head described in Sec. 2. Subsequently, all baselines, including the Pick-and-Place baselines, were fully fine-tuned on the same set of downstream demonstrations, following the evaluation protocol outlined in [1], to ensure a fair comparison.
@@ 3.2 Franka Kitchen
Overview. Franka Kitchen [54] is a well-established benchmark for evaluating the efficacy of visual representations in facilitating the learning of visuomotor control policies from limited demonstrations. This benchmark comprises five distinct visuomotor control tasks, as shown in Fig. 8, each captured from two camera viewpoints.
@@ 3.3 Referring Expression Grounding
Overview. The goal of this task is to predict a bounding box of an object in a cluttered scene based on the nature language expression. This task offers the evaluation of language-conditioned scene understanding and object recognising ability, which serves as an important prerequisite for language-based robot manipulation. The benchmark is based on OCID-Ref Dataset [62], which provides representatives scenes in robotics settings. The benchmark also provides splits based on the clutter level.
@@
Evaluation details. We use a shallow MLP as detection head to regress the bounding box directly. We use the evaluation codebase provided by [7]. Similar to Franka Kitchen, we report the results stated in the paper for each baseline [7, 40, 42, 43, 45]. The evaluation metrics are the average precision at 0.25 IoU under each clutter level.
@@ 3.4 Real robot experiments
Evaluation details. Fig. 10 shows our real-robot environment and objects. We train both our model and CLIPort [1] on our manually collected training demos. We utilise a 6-degree-of-freedom (6-DoF) UR5 robotic arm, Robotiq 2F-85 two-finger gripper, and an Intel RealSense RGB-D camera for our real-world experiments. We capture the top-down RGB observation which covers the workspace of 60 cm $\times$ 30 cm, and the image from the camera is resized to 320 $\times$ 160 pixels.
Task details. Here we show the language template, variable factors and the success condition for each real robot task in Tab. 4. Images for each task are presented in the main paper.
@@
Pressing button Press the {color} button Button color, button location Correct button touched
@@ 4 Additional details
Figure 12: Examples of masked image prediction. We show the qualitative examples of our masked auto-encoders on the validation set. The text descriptions are shown below each sample.
@@
Dataset details. We build our Omni-Object Pick-and-Place (OOPP) dataset upon the previous benchmarsk [1, 9] in the PyBullet Gym enviornment [11]. We manually selected 180 real-scanned object classes from the OmniObject3D dataset [10], focusing on those suitable for robotic manipulation, resulting in a total of 3,200 distinct instances. For each object, we reduced the mesh resolution to 20K faces to enable efficient rendering in simulation. Additionally, we filtered out objects that were either too large or too small (i.e., with dimensions less than 4cm or greater than 40cm). Our dataset includes full robot manipulation episodes, comprising of annotated robot actions, language instructions, and simulation-generated rewards. We use 160 object classes for training. For evaluating intra-class generalisation, we hold out a subset of instances from 20 categories included in the training set. For inter-class generalisation, we reserve 20 object categories that are entirely unseen during training. Fig. 13 illustrates the first 50 examples across different object classes, while Fig. 14 highlights examples of intra-class variation. The held-out categories for inter-class generalisation are sampled from four high-level semantic groups, as detailed in Table 5.
Figure 13: Examples of objects from different classes.
Figure 14: Examples of intra-class variance.
Table 5: Semantic group division of seen and unseen classes. We divided all objects into four semantic groups. We selected unseen classes from each group in proportion to the number of object classes it contains, ensuring an even distribution of unseen categories.
@@
dumpling, mooncake, rice_cake, puff, water_chestnut, mushroom,
@@
peach, apple, pear, potato, zongzi, pomegranate, onion, egg,
@@
belt, dish, flash_light, canned_beverage, fork, cup, teapot, book,
glasses_case, bowl, tape_measure, speaker, laundry_detergent,
@@
picnic_basket, hammer, shoe, hat, laptop, vase, ornaments, spanner
@@
toy_boat, toy_plant, toy_car, toy_plane, timer, whistle, doll,
table_tennis_bat, toy_motorcycle, drum, remote_control,
garage_kit, china, chess, Chinese_chess, rubik_cube, dinosaur
@@
hairpin, lotus_root, house (model), plant, dumbbell, package,